OpenStudy (anonymous):

simplify. -5 1/3 + (-3 9/18)

OpenStudy (anonymous):

-53/6

OpenStudy (anonymous):

1/3 = 6/18 -5 6/18 = -96 / 18 - 3 9/18 = - 63/18 adding : - 159 / 18 = -53 / 6